OpenStudy (anonymous):

Alli withdrew some money from the bank for shopping. she spent two thirds of what she withdrew on groceries. She spent $25 on a sweater. She spent half of what remained on a necklace. She went home with $15. How much did Alli withdraw from the bank.

OpenStudy (anonymous):

let the money withdrew = x \[(2/3)x + 25 + 15 + ((x - ((2/3)x + 25))/2) = x\] \[(2/3)x + 25 + 15 + (((1/3)x + 25)/2) = x\] \[(4/3)x + 80 + (1/3)x + 25 = 2x\] \[(5/3)x + 105 = 2x\] \[105 = 2x - (5/3)x\] \[105 = (6x - 5x)/3\] \[105 * 3 = x\]

OpenStudy (anonymous):

@aum Can you help in this?

OpenStudy (aum):

Let x be the money Ali withdrew. "she spent two thirds of what she withdrew on groceries." If she spent 2/3 * x on groceries she will have x - 2/3x = 1/3 * x = x/3 left. "She spent $25 on a sweater" So she will have x/3 - 25 left. " She spent half of what remained on a necklace" If she spent half she will have half left: 1/2 * (x/3 - 25) "She went home with $15" 1/2 * (x/3 - 25) = 15 multiply both sides by 2: (x/3 - 25) = 30 add 25 to both sides: x/3 = 55 multiply both sides by 3: x = 165. Ali withdrew $165.

OpenStudy (anonymous):

ok thats what i got the second time i did it thank you

OpenStudy (aum):

You are welcome.